OpenStudy (anonymous):

can anyone suggest me with the graph of f(x)= |x| + |x-1| + |x-2| +|x-3|

OpenStudy (rational):

for x >=3, the graph is a straight line : 4x - 6 for x <=0, the graph is a straight line : -4x + 6 need to study for x in between (0, 3)

OpenStudy (anonymous):

yeah that what I need, I remember someone taught me that graph... and need its minimum value and that x too

OpenStudy (sepeario):

used graphing tool on my mac :)

OpenStudy (anonymous):

we are not allowed to use electronic devices in exam :P @Sepeario

OpenStudy (sepeario):

Ok I would start with the fact that y must be at least 3.

OpenStudy (anonymous):

@rational what do you say.. define it in every interval then plot its graph?

OpenStudy (rational):

I think the minimum values for `|x| + x-1|+|x-2|+...` is achieved when x is the median of `0, 1, 2, ... `

OpenStudy (rational):

for example, the median of `0, 1, 2, 3, 4` is 2 therefore the minimum value of `|x|+|x-1|+|x-2|+|x-3|+|x-4|` is achieved when `x=2`

OpenStudy (anonymous):

but can this be generalized as for x E [1.2] the value is constant 4

OpenStudy (rational):

another example, the median of `2, 5, 6, 10, 12` is 6 therefore the minimum value of `|x-2|+|x-5|+|x-6|+|x-10|+|x-12|` must be achieved at `x=6`

OpenStudy (rational):

I think you will get a range of "x" for minimum when the number of terms is even

OpenStudy (anonymous):

but if the numbers are not close? then for the minimum value the biggest one have to be zero

OpenStudy (rational):

when the number of terms is odd, there will be a exact median, so u get only one "x" where the minimum is achieved

OpenStudy (rational):

I think it doesn't matter, the minimum is always achieved for the median value of x

OpenStudy (rational):

can you provide an example of where you think it could fail ?

OpenStudy (rational):

another example, the median of `2, 5, 6, 10, 10000000` is 6 therefore the minimum value of `|x-2|+|x-5|+|x-6|+|x-10|+|x-10000000|` must be achieved at `x=6`

OpenStudy (anonymous):

yeah you are correct..

OpenStudy (anonymous):

thanks :) @rational

OpenStudy (rational):

np

OpenStudy (alekos):

.